How many yards of material from a 24-yard length of cloth remain after 3 each 3 1/2 yards long, and 5 pieces each 2 1/4 yards long are removed?
sol:Total length of the cloth=24 yards
the length of the cloth after removing 3 pieces of length 31/2 yards=24-(3*3.5)
(i.e 1/2=0.5,3 1/2=3+0.5=3.5)
=24-10.5
=24.0-10.5=13.5
now the length of the cloth=13.5 yards
again the length of the cloth after removing 5 pieces of length 2 1/4 yards=13.5-(5*2.25)
(i.e;1/4=0.25,2 1/4=2+.25=2.25)
=13.5-11.25
=13.50-11.25
=2.25 yards
the length of the remaining cloth=2.25 yard

Answer by anantha(86)

sol:Total length of the cloth=24 yards
the length of the cloth after removing 3 pieces of length 31/2 yards=24-(3*3.5)
(i.e 1/2=0.5,3 1/2=3+0.5=3.5)
=24-10.5
=24.0-10.5=13.5
now the length of the cloth=13.5 yards
again the length of the cloth after removing 5 pieces of length 2 1/4 yards=13.5-(5*2.25)
(i.e;1/4=0.25,2 1/4=2+.25=2.25)
=13.5-11.25
=13.50-11.25
=2.25 yards
the length of the remaining cloth=2.25 yard